### v1

# Bridging the Gap Between Saliency Prediction and Image Quality Assessment thanks: This study was supported by Russian Science Foundation under grant 24-21-00172, https://rscf.ru/en/project/24-21-00172/

Abstract
Over the past few years, deep neural models have made considerable advances in image quality assessment (IQA). However, the underlying reasons for their success remain unclear, owing to the complex nature of deep neural networks. IQA aims to describe how the human visual system (HVS) works and to create its efficient approximations. On the other hand, Saliency Prediction task aims to emulate HVS via determining areas of visual interest. Thus, we believe that saliency plays a crucial role in human perception.
In this work, we conduct an empirical study that reveals the relation between IQA and Saliency Prediction tasks, demonstrating that the former incorporates knowledge of the latter. Moreover, we introduce a novel SACID dataset of saliency-aware compressed images and conduct a large-scale comparison of classic and neural-based IQA methods. Supplementary code and data will be available at https://github.com/Alexkkir/SACID.
Index Terms:
Image Quality Assessment, Visual Saliency Prediction, Explainable AI
I Introduction
Image Quality Assessment (IQA) aims to estimate perceptual image quality similar to the human visual system (HVS). It can be useful in enhancing the visual experience of humans in various applications such as image compression, restoration, acquisition, editing, and generation.
Despite IQA models achieving performance similar to human assessors on various benchmarks, the inner mechanisms of IQA models remain unclear. A question of interest is whether these IQA models implicitly incorporate characteristics from the human visual system (HVS), especially in focusing attention on certain image regions, a principle known as saliency [1, 2]. Saliency Prediction (SP), as a specialized field, has seen the emergence of SOTA models capable of determining the points of human visual interest.
Figure 1: Saliency and GradCAM identify important regions of image. Upper row – saliency maps, predicted with TranSalNet[3]. Bottom row – GradCAMs, build for Baseline-EfficientB0.
In this work, we explore the connection of IQA and SP. Our main contributions are as follows:
  • We propose a methodology to extract saliency maps from trained IQA models, which reveals that learning-based IQA methods incorporate saliency in their predictions and can outperform strong SP baselines like center-prior.
  • We present a method for parameter-free dual-task training strategy for IQA and Saliency Prediction, which reveals that these tasks are connected and can be solved simultaneously without quality drops.
  • We conduct a subjective study with 1400+ assessors to evaluate the effectiveness of existing IQA metrics for non-uniform compressed content employing various saliency-aware coding strategies.
II Related Work
II-A Image Quality Assessment
IQA aims to measure the perceptual quality of an image by assessing the extent of image artifacts and distortions. According to the requirement for ground-truth reference images, most IQA techniques can be categorized as Full-Reference (FR) or No-Reference (NR). The latter scenario is more relevant for real-world applications [4].
Early approaches based on natural scene statistics (NSS) [5, 6, 7, 8, 9], employed hand-crafted features and a regression model above them to predict the mean opinion score (MOS). However, as deep learning has evolved, neural network-based approaches have become popular. Usually, they consist of a convolutional or transformer backbone, trained end-to-end with a regression head. Many studies have suggested iterative improvements of these architectures. HyperIQA [10] proposed a multiscale feature; MUSIQ [11] employed multiscale inputs and a vision transformer backbone to process images with varying sizes and aspect ratios.
Several authors [12, 13] have proposed methods for creating paired models that can serve in both FR and NR situations. For example, TOPIQ utilizes features from the original, distorted, and difference images at each layer.
II-B Saliency in Image Quality Assessment
Prior research indicates a connection between saliency and Image Quality Assessment [12, 14, 15, 16, 17, 18, 19, 20, 21 ]. Efforts are underway to enhance IQA-model performance by incorporating saliency [15, 22, 23, 24, 25, 26, 27, 28]. Early studies [9] have demonstrated improvement in NSS-based metrics like PSNR and SSIM by reweighting the error map based on saliency. In SGDNet [22], authors introduced a separate head to predict saliency and then applied its predictions to reweight features from the final layer.
TranSLA [24] incorporates saliency as a query branch in cross-attention, while HVS-5M [23] employs a pre-trained SP model to reweight features from the final layer. In SCVC [26], authors split the image into patches and aggregated their scores using a Gaussian function and saliency maps (SM). Furthermore, according to [12], LPIPS [29] performance can be enhanced by spatially reweighting feature maps. However, several studies [14] report that modern video quality assessment models (VQA) face only slight improvement from using saliency. Notably, our research reveals that models already possess knowledge of saliency, enabling them to incorporate it into their predictions.
II-C AI Explainability
Even though neural networks reached a human level performance in numerous tasks, they remain a black box. Special methods thus have been developed to provide insight into their internal operations. The first effort for explaining CNNs, CAM [30], considers a simple case of a convolutional network with a head consisting of a single linear layer after global average pooling (GAP). To obtain attention maps, features from the last layer are summed with weights from the fully connected layer. GradCAM [31] extends CAM to models with arbitrary head, employing gradients w.r.t. the model’s predictions rather than relying on weights. Subsequent works [32, 33, 34] propose heuristics for improving the method.
In another vein, some researchers have studied benchmarking explanation maps. In [32] they suggest masking the most important image areas and monitoring changes in model confidence. The authors of [35] retrained the model on perturbed images to reduce domain shift. Work [36] introduced a metric, area over the perturbation curve, as a measure of explanation-map fidelity. In [37] the researchers propose a novel method for masking image areas.
Furthermore, several works address explainability in IQA. For instance, in [13] authors showed that IQA models only require half of an image to make accurate predictions. To identify essential areas, they divided each image into 12 square patches and examined predictions of the transformer-based IQA model as they masked their various combinations. Then they measured the accuracy of SOTA IQA model on trivial and important regions that make up half of the area of the original image. They found out that when using the former, the quality decreases, while when using the latter, the quality of the model is almost preserved. We simplify this approach and show that masking pixels based on saliency maps or GradCAM yields comparable results.
III Experiments
III-A Extracting Saliency from IQA models
TABLE I: Performance comparison on SALICON[38] dataset. GradCAM-extracted saliency maps from IQA models outperform center-prior baseline, but are surpassed by SOTA SP methods. The best results are bold, the second best are underlined, and the third best are italics.
Type Method NSS $\uparrow$ SIM $\uparrow$ CC $\uparrow$ KLD $\downarrow$
Dummy CenterPrior 0,582 0,534 0,541 0,784
GradCAM from IQA Baseline-EfficientNet 0,604 0,584 0,637 0,700
Baseline-ResNet50 0,621 0,555 0,590 0,782
TOPIQ [12] 0,597 0,552 0,572 0,759
DBCNN [39] 0,579 0,566 0,590 0,707
CLIP-IQA+ [40] 0,584 0,550 0,589 0,741
SOTA SP MSINet [41] 0,612 0,767 0,891 0,252
TranSalNet [3] 0,613 0,776 0,903 0,258
Figure 2: Masking input images according to saliency and GradCAM maps with black (upper row) and ImageNet mean (bottom row). We note a strong relation between the behavior of correlations when masking with saliency and GradCAM maps.
We started our experiments by testing the hypothesis that ground-truth saliency (e.g. from eye-tracker) correlates with explanation GradCAM maps of IQA models. In our experiments we used TOPIQ [12], DBCNN [39] and CLIP-IQA [40]. We also trained our simple baseline model, consisting of a backbone, GAP pooling, and MLP head. We tested two backbones: EfficientNet-B0 [42] and ResNet-50 [43]. To build GradCAM maps, we choose the HiResCAM [33] method and applied smoothing through augmentations and SVD decompositions of feature maps as recommended in [44].
We used straightforward saliency baselines – CenterPrior, a Gaussian distribution that approximates the saliency averaged over the dataset – and two SOTA saliency prediction models [41, 3]. We evaluate the model quality based on common saliency metrics – NSS, SIM, CC, and KLD calculated on validation split of SALICON dataset [38]. Before calculating metrics, we employed map transforms [45].
Results are presented in Tab. I. GradCAM better predicts saliency than center-prior, indicating that IQA models understand saliency distribution and allow extracting such proxy SM in a zero-shot mode.
III-B Saliency Masking
Inspired by [13], we employed saliency and GradCAM maps for masking image regions. For better interpretability, we masked different image portions with two strategies: Most Relevant First (MoRF) and Least Relevant First (LeRF) [36]. We applied perturbations such as filling with black color and ImageNet mean. Before thresholding explanation maps, we apply Gaussian blur with a large kernel, to ensure we can calculate all quantiles. Fig. 1 shows examples of images after applying the first perturbation. We evaluated the correlation drop on the KonIQ-10k dataset; Fig. 2 presents the results.
Masking of salient areas decreases quality more than masking of trivial areas does. The same observation holds for GradCAM. Interestingly, masking with saliency achieves a larger performance gap between LeRF and MoRF regions compared to [13], even though the latter employs knowledge of pre-trained IQA models. This finding indicates that saliency addresses the regions that are crucial for IQA models. Furthermore, we note that the behavior of models when masked with GradCAM maps and saliency exhibits significant similarity, which we can interpret as an existing relationship between the attention of IQA models and humans.
III-C Dual-Task Training
After discovering a close connection between IQA and saliency prediction (SP), we decided to train a model on both tasks simultaneously and determine whether we could do so without decreasing performance. Our effort took two approaches.
In our first approach, denoted as Baseline-Sal-Loss, we simply added to our baseline model a small decoder, consisting of a single convolution with a 1×1 kernel followed by a sigmoid function. Our second approach employed the model’s GradCAM as the second output by incorporating it into the loss function. This enabled us to avoid adding any extra parameters to the model. The GradCAM calculation is as follows:
$L^{C}_{\text{GradCAM}}=\operatorname{ReLU}\underbrace{\left(\sum_{k}\alpha^{c}_{k}A^{k}\right)}_{\mathclap{\text{linear imputation}}},\quad\alpha^{c}_{k}=\overbrace{\frac{1}{Z}\sum_{i}\sum_{j}}^{\mathclap{\text{global average pooling}}}\underbrace{\frac{\partial y^{c}}{\partial A^{k}_{ij}}}_{\mathllap{\text{gradients via backprop}}}$
In our case, $c=1$. To ensure differentiability, we detached $\alpha^{c}_{l}$ from the computational graphs and calculated gradients only over channel maps. This resulted in a model called Baseline-GradCAM-Loss.
We used a random 80% of the KonIQ-10k dataset [46] for training and the remaining 20% for testing as well as CLIVE [47], SALICON [38], and CAT2000 [48] datasets. As KonIQ-10k doesn’t provide SM, we used proxy SM generated with MSINet[41]. We repeated each experiment 10 times and presented averaged results in Tab. II. We compared these dual-task models with the baseline model trained solely on IQA (with III-A saliency extraction technique), CenterPrior, and SOTA saliency models. The findings indicate that both approaches allow us to train models on both tasks without decreasing the performance of IQA while significantly improving SP results.
III-D Non-uniform-Compression Dataset
TABLE II: Simultaneous learning of two tasks.
Method IQA Saliency Prediction
KonIQ-10k[46] CLIVE[47] SALICON[38] CAT2000[48]
SROCC $\uparrow$ PLCC $\uparrow$ SROCC $\uparrow$ PLCC $\uparrow$ NSS $\uparrow$ SIM $\uparrow$ CC $\uparrow$ KLD $\downarrow$ NSS $\uparrow$ SIM $\uparrow$ CC $\uparrow$ KLD $\downarrow$
Center Prior - - - - 0.578 0.536 0.544 0.779 0.303 0.599 0.771 0.672
Baseline 0.913 0.931 0.862 0.848 0.604 0.584 0.637 0.700 0.308 0.613 0.781 0.585
Baseline-Sal-Loss 0.914 0.930 0.845 0.839 0.618 0.688 0.803 0.429 0.372 0.665 0.818 0.433
Baseline-GradCAM-Loss 0.912 0.931 0.852 0.840 0.615 0.703 0.825 0.396 0.382 0.676 0.826 0.403
MSINet[41] - - - - 0.612 0.767 0.891 0.252 0.370 0.664 0.820 0.421
TranSalNet[3] - - - - 0.613 0.776 0.903 0.258 0.369 0.670 0.829 0.407
[…]
Subsequently, we calculated SROCC, PLCC and KRCC for each metric from the PYIQA toolbox [50], saliency-aware versions of PSNR and SSIM (EW-PSNR and EW-SSIM [9]), our model variants (Baseline, Baseline-Sal-Loss, and Baseline-GradCAM-Loss), and salient deep-learning models (SGDNet[22], etc.). Each image was provided in its original resolution to maximize the performance of most models. KRCC was calculated in groups with different both sequences and bitrates; thus we denote it as $\text{KRCC}^{\dagger}$.
Results are listed in Tab. III. Notably, salient versions of PSNR and SSIM remarkably outperformed the originals. But Baseline, Baseline-sal-loss, and Baseline-gradcam-loss models demonstrated similar performance, implying that saliency fails to enhance deep-learning metrics significantly. We attribute this to the reliance of methods on conventional datasets, where most distortions are uniform, despite being trained with saliency. Thus, IQA models are limited by current compression standards and may show lower correlations in the non-uniform compression domain.
We also observed intriguing disparities regarding to TOPIQ model: TOPIQ-NR achieved notably higher scores than TOPIQ-FR. Moreover, most NR metrics are ranked above FR (as depicted in Fig. 3).
IV Discussion
TABLE III: Quantitative results on SACID.
Type Method SROCC $\uparrow$ PLCC $\uparrow$ $\text{KRCC}^{\dagger}$uparrow
NSS based MS-SSIM [6] 0.807 0.849 0.013
BRISQUE [8] 0.817 0.833 0.144
PSNR 0.822 0.858 0.009
SSIM [5] 0.835 0.880 0.004
VIF [7] 0.848 0.891 0.052
EW-SSIM [9] 0.850 0.867 0.272
EW-PSNR [9] 0.875 0.893 0.356
FR AHIQ [51] 0.789 0.817 0.091
LPIPS [29] 0.818 0.854 0.024
PieAPP [52] 0.833 0.869 0.069
TOPIQ-FR [12] 0.836 0.848 0.257
DISTS [53] 0.871 0.904 0.178
NR MANIQA [54] 0.752 0.773 0.128
TReS [55] 0.765 0.800 0.117
HyperIQA [10] 0.784 0.802 0.124
SGDNet-None [22] 0.820 0.853 0.302
PaQ2PiQ [56] 0.848 0.858 0.237
SGDNet-Output [22] 0.854 0.879 0.374
Baseline-GradCAM-Loss 0.863 0.865 0.248
CLIP-IQA+ [40] 0.869 0.896 0.196
Baseline 0.871 0.881 0.217
TOPIQ-NR [12] 0.876 0.886 0.257
Baseline-Sal-Loss 0.880 0.886 0.259
MUSIQ [11] 0.888 0.899 0.324
DBCNN [39] 0.901 0.917 0.331
Initially, our experiments focused on developing a saliency-aware IQA metric and exploring its potential to enhance IQA performance, given their evident relationship. However, we encountered challenges, prompting a shift in focus to investigate the underlying reasons.
We visualized GradCAMs of several IQA-model and noted their similarity to SM even though the IQA should pay more attention to areas with artifacts. Hence, we hypothesized that learnable models inherently understand saliency patterns. To confirm this, we introduced a method for extracting saliency maps from IQA models, showed that they mostly rely on salient regions, and proved the possibility of training a single model without additional parameters for both tasks without performance degradation.
Figure 3: Correlations of top-10 full-reference and no-referene methods on proposed SACID dataset.
We also questioned the adequacy of existing datasets, as they primarily contain images compressed with conventional codecs. To overcome this limitation, we created a new dataset consisting of nonuniformly compressed images. We tested various IQA metrics on this dataset and confirmed that incorporating saliency can enhance simple metrics such as PSNR and SSIM, however, deep-learning-based failed to see a similar boost in quality. Additionally, we observed that no-reference metrics outperformed full-reference metrics.
As a future work, we are considering improving the robustness of metrics to adversarial attacks by utilizing saliency.
V Conclusion
This study examined the relationship between saliency prediction and image-quality assessment. We propose a technique for extraction saliency maps from IQA models and empirically demonstrate their reliance on salient regions. Additionally, we propose a parameter-free dual-task learning approach without sacrificing quality. Finally, we curated a dataset of non-uniformly compressed images and performed large-scale comparison, which reveals that NR metrics are better suited for saliency-aware compression domain. We conclude that saliency can remarkably improve NSS-based methods, while learning-based methods face slight improvement.